OpenStudy (anonymous):

Point X is located at (2, -6) and point Z is located at (0, 5). Find the point Y that is located 1/5 the distance from point X to point Z. a. (1.4, -3.6) b. (1.6, -3.8) c. (1.8, -4) d. (2, -4.2) PLEASE HELPPP! I WILL GIVE YOU A MEDAL IF YOU'RE CORRECT:)

OpenStudy (debbieg):

I think you just have to look at this as a variation of midpoint. But instead of being "midway" between the points, you want to be 1/5 of the way FROM x TO z. I would probably first notice that 1/5 = 0.2. Then for the x's, you are going from 2 toward 0, and you want to go 0.2 of that distance. So that would be 2 - 0.2(distance between 2 and 0) = 2 - 0.2(2).... that gives you your x coordinate. Similarly for the y's, you are going from -6 toward 5, and you want to go 0.2 of that distance. So that would be -6 + 0.2(distance between -6 and 5) = -6 + 0.2(11).... that gives you your y coordinate.

OpenStudy (debbieg):

Does that make sense?

OpenStudy (anonymous):

whats the answer cause i got an answer that wasnt listen what is the correct answer so i know what i am looking to find? pleasee

OpenStudy (debbieg):

Did you understand what I wrote above? I showed you exactly what to do: The x-coordinate= 2 - 0.2(distance between 2 and 0) = 2 - 0.2(2) The y-coordinate = -6 + 0.2(distance between -6 and 5) = -6 + 0.2(11) Do you see why? If you compute them both like that, the answer is there. :)

OpenStudy (anonymous):

I did! but i got 11 for the second one?

OpenStudy (debbieg):

-6 + 0.2(11) = -6 + 2.2 = ?

OpenStudy (anonymous):

oh thanks i got it!:) could you help me with another one?

OpenStudy (anonymous):

Using the image below, find the point that divides the line segment CD into a ratio of 4:1. a. (5.8, 2) b. (6, 6.2) c. (6.5, 2.8) d. (7, 3.2)

OpenStudy (debbieg):

hold on a sec....

OpenStudy (debbieg):

OK, I wanted to check my answer. I think it's the same idea as the one we just did. the only difference is that they don't tell you if you want the point closest to (5,1) or the point closest to (9,6) (because you can divide the line segment in that ratio at either end). But by saying a ratio of 4:1, that really means into 5 equal "parts", right? So one part of the line has 4 parts and the other has 1 part. So, again, this is 1/5 of the distance between the points. Points are (5,1) and (9,6), so first trying with the "short end" near the (5,1), you are going 0.2 of the distance from (5,1) to (9,6). So the computations are just as above: x-coordinate = 5 + 0.2(distance between 5 and 9) y=coordinate = 1 + 0.2(distance between 1 and 6)

OpenStudy (debbieg):

I bet you can get it from that. :)

OpenStudy (anonymous):

i got 5.2 and 1.2 and it was none of the answers iabove

OpenStudy (debbieg):

Kind of strange how the question is worded, since like I said, there are TWO points that divide the line segment into that same ratio - the other is near the other end. You would find it the same way as above, except using 0.8 where we used 0.2... but when you do that, you don't get those coordinates as a choice.

OpenStudy (debbieg):

can you show me your work? What exactly did you evaluate...?

OpenStudy (debbieg):

Show me how you calculated these: x-coordinate = 5 + 0.2(distance between 5 and 9) y=coordinate = 1 + 0.2(distance between 1 and 6)

OpenStudy (debbieg):

oh, I see what you did... you just added 0.2 to each of the starting points.

OpenStudy (debbieg):

it is 0.2 * (distance between coordinates) You have to take 0.2 of the distance between the two x coordinates, and add that to 5. You have to take 0.2 of the distance between the two y coordinates, and add that to 1.

OpenStudy (anonymous):

i got b is what correct?

OpenStudy (debbieg):

No. What are you using for the distances? What is the distance between 5 and 9? What is the distance between 1 and 6?

OpenStudy (debbieg):

Notice that 6.2 certainly can't be your y-coordinate BETWEEN those two points, it is larger than the larger of the two y's.

OpenStudy (anonymous):

oh it was d thank you!!

OpenStudy (debbieg):

What are you using for the distances? What is the distance between 5 and 9? What is the distance between 1 and 6?